# Assessment of Simulation-based Inference Methods for Stochastic Compartmental Models

### v2

@@ 5.3 Reducing Parameter Dependencies Improves Posterior Alignment
Complementing the analysis of the full SEIR2v model, we evaluated the reparametrized version of the SEIR model to explore how reduced parameter dependencies affect inference results for PF and CNF.
Posterior distributions from CNF and PF showed close alignment across parameters, indicating similar exploration of the parameter space (Figure 5A). Both methods yielded comparable MAP estimates near the true values, with minor deviations for $e_{0}$ and $I_{0}$, where PF shifted toward lower values. Epidemic trajectories fit the data equally well (Figure 5B) and RMSE values were similar across CNF, PF, and the true parameters, with PF showing marginally better agreement on the infection count data (Table S21).
Additionally, we evaluated calibration and convergence diagnostics. CNF showed SBC miscalibration in ECDF plots (Figure S22), consistent with the results reported for the identifiable SIR model. PF, in contrast, achieved ESS exceeding 2000 (Supplementary Table S22), indicating well‑mixed chains and thorough posterior exploration.
@@
Table 6: MAP and RMSE for reparametrized SEIR model and dense datasets.
$\mathbf{r0}$ $\mathbf{e0}$ $\mathbf{s0}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
True Par. 1.36 22.00 51.00 150.00 500.00 0.0200 0.0245
CNF 1.34 21.33 50.27 145.46 489.71 0.0216 0.0236
PF 1.34 20.57 51.21 156.94 467.30 0.0319 0.0271
@@ 5.4 CNF and PF yield consistent results on real-world datasets
To complement the synthetic experiments, we applied both methods to data from a longitudinal cohort study of the COVID‑19 outbreak in Ethiopia [ethiopian_data]. Our setup differed from the original study in that we used a stochastic rather than deterministic formulation, offering greater flexibility and better reflecting the inherent randomness in infection dynamics. Due to the unavailability of ground truth, we used the published mean of the parameter samples as a reference.
Posterior approximations from CNF and PF shared a similar overall shape, with the PF posterior largely nested within the broader CNF and exhibiting reduced variance, particularly for the rate parameters $\beta$, $\kappa^{-1}$, and $\gamma^{-1}$ (Figure 6A). Trajectory ensembles derived from the inferred posteriors fit the data well, even under the relatively high noise level reported in the original study; one seroprevalence point was not captured by PF within its uncertainty intervals (Figure 6B). Despite this discrepancy, the PF MAP achieved RMSE values that matched or improved upon those reported in the original study. Although the CNF MAP deviated more strongly from both the published parameters and PF, the RMSE values of the respective distributions (Table S24) were comparable and showed improved accuracy for infection count data.
Diagnostics were consistent with patterns observed in controlled experiments: CNF exhibited good calibration alongside imperfect parameter recovery (Figure S21, Figure S25), while PF demonstrated convergence despite low effective sample sizes (Table S25, Table S26).
@@
Table 7: MAP and RMSE for full SEIR model for real dataset.
$\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa^{-1}}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I_{0}}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
Pub. Par. 16.70 5.00 0.0800 2.30 184.50 635.00 0.0737 0.0702
CNF 14.80 4.14 0.0882 3.16 183.04 550.16 0.1251 0.0928
PF 15.03 4.67 0.0858 3.36 160.35 752.08 0.0856 0.0606
@@ 6 Discussion
Responses to public health crises, such as outbreaks of infectious diseases, creates a need for informed decision making by realistically modeling the dynamics using stochastic models. Yet, fitting such models to observed data present many difficulties and several methods were developed address this challenge. In this work, we systematically evaluated two state-of-the-art methods for Bayesian parameter inference, Conditional Normalizing Flows (CNF) and pseudo-marginal particle MCMC, referred to as Particle Filters (PF), across a range of stochastic compartmental models. Our analysis encompassed synthetic and real-world data, varied model structure, and different levels of data sparsity, yielding crucial insights into the strengths and limitations of each methodology when applied in epidemiological modeling scenarios. Across all investigated scenarios, both methods produced accurate fits and informative posteriors. However, this study revealed systematic differences of uncertainty in the parameter space due to different posterior geometries and model design choices.
@@
In the context of an identifiable stochastic model, such as the SIR model, both CNF and PF yielded highly consistent posterior distributions and closely matched predictive epidemic trajectories. The marginal and joint posterior estimates and MAP estimates were near the data-generating parameters and the root mean square error (RMSE) values confirmed high predictive accuracy for both methods, despite imperfect calibration of the CNFs. Additionally, CNF exhibited very good recovery plots in this setting. As a exact method, PF showed a good recovery of the true parameters in this setting, that were closely matched by the CNF. Together, this indicates that, once trained, CNF can deliver inference quality on a par with an MCMC based baseline while enabling possible amortization across different datasets and yielding high computational efficiency in the inference phase.
@@
Expanding the analysis, we applied both methods to a more complex stochastic two-variant SEIR model with practical non-identifiabilities and parameter couplings. In this setting, both methods fit the data well and achieved similar predictive power and uncertainty, yet PF produced narrower marginal posteriors that were largely enclosed by the broader CNF posteriors and simultaneously showed low effective sample sizes for several parameters. This pattern is consistent with resampling in PF concentrating particles in high-likelihood basins, i.e., on the fit to the data. This property can limit tail exploration in the parameter space especially when the posterior surface contains ridges or mountain-pass structures. CNF achieved good calibration results in this setting, although its recovery plots indicated that parameter recovery was not perfect. The wider posteriors for CNF contrasts with prior literature where amortized inference often underestimates uncertainty [hermans2022trust]. Despite these differences, both methods captured the inference data accurately and showed strong predictive performance, as evidenced by low and similar RMSEs. Their application in this setting shows that amortized methods may struggle with calibration, yet still provide strong predictive performance and exact MCMC based method estimate the bulk of the posterior well while underestimating its width caused by too less parameter space explorance.
@@
Using sparsely sampled datasets with different missingness patterns did not change the qualitative estimation results of both methods. But they inherently treat sparse datasets differently. While CNF methods can accommodate such dataset variability, amortizing over all relevant combinations of time points may become computationally expensive and risks unreliable extrapolation if not carefully designed. This challenge is specific to setups that aim for broad amortization; alternatively, one can train CNFs for fixed missingness patterns. In contrast, PF methods naturally handle non-equidistant and partially missing data via their sequential simulation and weighting scheme, which decomposes the likelihood over time points and observables without requiring amortization or any other adaptation of the method.
@@
Finally, applying both methods to a real dataset from a longitudinal cohort study in Ethiopia confirmed the simulation study findings and yielded broadly similar posterior approximations and epidemic trajectories. Although obtained MAP estimates differed from previously published parameter values, especially for CNFs, both methods captured the data dynamics and produced model fits with RMSEs close to or improving upon the deterministic analysis [ethiopian_data]. By demonstrating their effectiveness in a real-world setting, our analysis underscores the potential of these methods to inform public health strategies and decision-making processes.
@@
We also compared the computational efficiency of the two methods. To ensure a meaningful comparison, both methods were calibrated to produce the same number of samples from the parameter posterior, with CNF’s neural network trained on a CPU. Despite this constraint, CNF is roughly 10 times faster than PF, due to PFs sequential nature and the higher number of simulator calls. The bulk of CNF’s computational time is spent on training, but once trained, inference can be performed in mere seconds. CNF’s amortization allows multiple inferences, providing, in this scenario, a significant advantage over PF, which requires separate computations for each run. To achieve greater speed, CNF can make use of a parallelized simulator or training the neural network on a GPU. Meanwhile, PF could benefit from the use of more advanced filtering algorithms such as guided PFs [snyder_2011, chopin_2020] or using correlated particles [wiqvist_2021] to avoid degeneracy or too high particle numbers. Switching to a more advanced outer MCMC sampler such as a Gibbs sampler [chopin2015] or a Metropolis-Adjusted Langevin algorithm [lowe2023accelerating] and parallelization over particles could speed-up the inference procedure as well. Additionally, the lack of amortization renders assessment of calibration and hyperparameter tuning for PF too computationally expensive.
@@
The scope of this work leads to several limitations. The observation model employed additive Gaussian noise for clarity and comparability; overdispersed counts or serological adjustments may be better represented by negative‑binomial or beta‑binomial formulations in other settings. Our use of an Euler–Maruyama integration with positivity clipping introduces discretization error whose interaction with inference can lead to numerical instabilities near boundaries or in small populations. Algorithmic choices were guided by the ease of their applicability in praxis: We focused on a bootstrap PF within PMMH and a tuned CNF architecture with consistent hyperparameters and fixed simulation budget across our datasets. Alternative PF proposals (e.g., guided/auxiliary filters or tempering within SMC) and alternative SBI estimators (likelihood or ratio estimation) or more exhaustive hyperparameter-tuning could shift efficiency–accuracy trade‑offs. To avoid out‑of‑distribution failure modes, CNFs were trained per experiment rather than globally amortized across all data regimes, which tempers their wall‑clock advantage in this study; broader amortization remains attractive but demands careful conditioning on meta‑data such as time‑grid irregularity and noise scales [radev2023bayesflow, wang2024missing].
@@
In the SBI literature, recent advances focus on workflows that narrow the gap between neural SBI and Monte Carlo by using learned proposals inside sequential Monte Carlo or MCMC [papamakarios2019sequential, hermans2020likelihood]. Variational and amortized SMC approaches train proposal families to reduce weight degeneracy and improve effective sample sizes; conversely, particle methods can provide high‑quality samples to bootstrap or regularize neural estimators. Looking forward, future work in this context of hybrid approaches could focus on a flow trained in the spirit of neural posterior estimation that proposes parameter values to a PMMH kernel or guide latent‑state proposals. These methods could combine the tail coverage and speed of neural estimators with the asymptotic correctness and model flexibility of pseudo‑marginal schemes. Recent generative modeling advancements, i.e., flow matching and diffusion‑style transports, offer further avenues for robust posterior learning with stronger inductive biases and improved stability [lipman2022flow, draxler2024free, arruda2025]. Additionally, the use of adaptive prior schemes where the prior is updated in light of the available data, through approaches such as empirical Bayes, hierarchical modeling, or iterative refinement, may help to mitigate issues of non-identifiability while improving robustness. Such strategies may enhance performance in particularly challenging applications that use real world data where information is both sparse and noisy.

### v3

@@ 5.4 Reducing Parameter Dependencies Improves Posterior Alignment
Complementing the analysis of the full SEIR2V model, we evaluated the reparametrized version of this model to explore how reduced parameter dependencies affect inference results for PF and CNF.To this end, we analysed twelve synthetic datasets in both the full and the reparametrized settings. In the full SEIR2V model, two datasets (r-2 and r-5) did not yield converged Markov chains under our PMMH workflow.
In the reparametrized setting, posterior distributions from CNF and PF showed close alignment across parameters, indicating similar exploration of the parameter space (Figure 6A). Unlike in the full SEIR2V model, the posterior approximations exhibited no signs of long‑tailed behaviour. This is reflected in the 1‑Wasserstein distances in log‑parameter space, which were numerically smaller under reparametrization and decreased more strongly than the within‑method baselines (CNF–CNF and PF–PF). Moreover, CNF–PF distances were of the same order as the variability observed across independent runs of each method, indicating markedly improved cross‑method alignment. Epidemic trajectories fit the data equally well (Figure 6B) with energy scores between CNF, PF, and the true parameters being close to each other in both model settings (Table 9, Table 10).
Additionally, we evaluated calibration and convergence diagnostics. While results in the full model setting were consistent with the ones presented above, CNF showed SBC miscalibration in ECDF plots in the reparametrized setting(Supplementary Fig. S6.14), aligning with the results reported for the identifiable SIR model and hinting at a potential offset in the posterior. PF, in contrast, achieved ESS exceeding 2000 (Supplementary Table S6.2), indicating well‑mixed chains and thorough posterior exploration.
@@
Table 9: Predictive Energy Score for the reparametrized SEIR2V model with dense data.
d-1-1 d-1-2 r-1 r-2 r-3 r-4
True Param. $0.1133$ $0.1153$ $0.0973$ $0.1386$ $0.0986$ $0.0997$
CNF 0.1098 $0.1066$ $0.0832$ $0.1176$ $0.1003$ $0.1035$
PF 0.1074 $0.1043$ $0.0799$ $0.1163$ $0.0935$ $0.1052$
r-5 r-6 r-7 r-8 r-9 r-10
True Param. $0.1455$ $0.1206$ $0.1463$ $0.1226$ $0.0913$ $0.0628$
CNF 0.1264 $0.1166$ $0.1418$ $0.0980$ $0.0818$ 0.058 030.058\,03
PF 0.1277 $0.1059$ $0.1384$ $0.0991$ $0.0792$ $0.0587$
Table 10: Predictive Energy Score for the full SEIR2V model with dense data from the reparametrized setting.
d-1-1 d-1-2 r-1 r-2 r-3 r-4
True Param. $0.1132$ $0.1148$ 0.097 540.097\,54 $0.1388$ 0.098 260.098\,26 0.099 690.099\,69
CNF $0.108$ $0.1067$ 0.092 690.092\,69 $0.1227$ $0.1004$ $0.1063$
PF $0.1062$ $0.1035$ 0.079 610.079\,61 0.094 380.094\,38 $0.1005$
r-5 r-6 r-7 r-8 r-9 r-10
True Param. $0.1456$ $0.1203$ $0.146$ $0.124$ 0.091 420.091\,42 0.063 120.063\,12
CNF $0.1327$ $0.1127$ $0.1445$ 0.098 980.098\,98 0.080 960.080\,96 0.057 730.057\,73
PF $0.1059$ $0.1382$ 0.098 920.098\,92 $0.0796$ 0.057 370.057\,37
@@
Full diagnostics, including Wasserstein distance analyses, calibration plots, convergence statistics, and parameter‑recovery results, are provided in Supplementary Information 6 (reparametrized model) and Supplementary Information 7 (full model with datasets for the reparametrized model).
@@ 5.5 CNF and PF yield consistent results on real-world datasets
To complement the synthetic experiments, we applied both methods to data from a longitudinal cohort study of the COVID‑19 outbreak in Ethiopia [Gudina et al., 2021]. Our setup differed from the original study in that we used a stochastic rather than deterministic formulation, offering greater flexibility and better reflecting the inherent randomness in infection dynamics. Due to the unavailability of ground truth, we used the published mean of the parameter samples as a reference.
Posterior approximations from CNF and PF shared a similar overall shape, with the PF posterior largely nested within the broader CNF and exhibiting reduced variance, particularly for the rate parameters $\beta$, $\kappa^{-1}$, and $\gamma^{-1}$ (Figure 7A). Trajectory ensembles derived from the inferred posteriors fit the data well, even under the relatively high noise level reported in the original study; one seroprevalence point was not captured by either method within their uncertainty intervals (Figure 7B). Predictive energy scores between CNF and PF improved on the energy score calculated from the the originally published parameter set, indicating that both inference methods provide a better probabilistic fit to the observed data than the originally published parameter set (Table 11).
Diagnostics were consistent with patterns observed in controlled experiments: CNF exhibited good calibration alongside imperfect parameter recovery, while PF demonstrated convergence despite low effective sample sizes.
@@
A Posterior approximations based on 10,000 samples, publication parameters and their reported standard deviation are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Table 11: Predictive Energy Score for the full SEIR2V model with real-world data.
Ethiopia
Pub. 0.131
CNF 0.107
PF 0.096
@@
Full diagnostics, including calibration plots, convergence statistics, and parameter‑recovery results, are provided in Supplementary Information S8.
@@ 6 Discussion
Responses to public health crises, such as outbreaks of infectious diseases, creates a need for informed decision making by realistically modeling the dynamics using stochastic models. Yet, fitting such models to observed data present many difficulties and several methods were developed to address this challenge. In this work, we systematically evaluated two state-of-the-art methods for Bayesian parameter inference, Conditional Normalizing Flows (CNF) and pseudo-marginal particle MCMC, referred to as Particle Filters (PF), across a range of stochastic compartmental models. Our analysis encompassed synthetic and real-world data, varied model structure, and different levels of data sparsity, yielding crucial insights into the strengths and limitations of each methodology when applied in epidemiological modeling scenarios. Where feasible, we calculated a reference posterior using Hamiltonian Monte Carlo (HMC) on a finely discretized version of the stochastic models. Across all investigated scenarios, both methods produced accurate fits and informative posteriors. However, this study revealed systematic differences of uncertainty in the parameter space due to different posterior geometries and model design choices.
@@
In the context of small stochastic models, such as the SIR model and the SIS model, both CNF and PF yielded highly consistent posterior distributions and closely matched predictive epidemic trajectories. The marginal and joint posterior estimates and MAP estimates were near the data-generating parameters and the predictive energy scores confirmed high predictive accuracy for both methods. CNFs exhibited miscalibration for the SIR model, which resulted in small offsets relative to PF and HMC. In contrast to that, CNF exhibited very good recovery plots in these settings. As a exact method, PF showed a good recovery of the true parameters and a high agreement with the reference posterior distributions, that were closely matched by the CNF. Together, this indicates that, once trained, CNF can deliver inference quality on a par with an exact MCMC based baseline while enabling possible amortization across different datasets and yielding high computational efficiency in the inference phase but miscalibration needs to be handled and analysed carefully.
@@
Expanding the analysis, we applied both methods to a more complex stochastic two-variant SEIR model with practical non-identifiabilities and parameter couplings. In this setting, both methods fit the data well and achieved similar predictive power and uncertainty, yet PF produced narrower marginal posteriors that were largely enclosed by the broader CNF posteriors and simultaneously showed low effective sample sizes for several parameters. This pattern is consistent with resampling in PF concentrating particles in high-likelihood basins, i.e., on the fit to the data. This property can limit tail exploration in the parameter space especially when the posterior surface contains ridges or mountain-pass structures. Additionally, PF based methods are, as most MCMC methods, highly dependent on the choice of the initial parameter value. Uninformative initial parameter values lead to degenerate likelihood estimate and therefore stuck Markov chains. CNF achieved good calibration results in this setting, although its recovery plots indicated that parameter recovery was not perfect. The wider posteriors for CNF contrasts with prior literature where amortized inference often underestimates uncertainty [Hermans et al., 2022]. Despite these differences, both methods captured the inference data accurately and showed strong predictive performance, as evidenced by low and similar predictive energy scores. Their application in this setting shows that amortized methods may struggle with calibration, yet still provide strong predictive performance and exact MCMC based method estimate the bulk of the posterior well while underestimating its width caused by too less parameter space explorance.
@@
Using sparsely sampled datasets with different missingness patterns did not change the qualitative estimation results of both methods. But the two methods inherently treat sparse datasets differently. While CNF methods can accommodate such dataset variability, amortizing over all relevant combinations of time points may become computationally expensive and risks unreliable extrapolation if not carefully designed. This challenge is specific to setups that aim for broad amortization; alternatively, one can train CNFs for fixed missingness patterns. In contrast, PF methods naturally handle non-equidistant and partially missing data via their sequential simulation and weighting scheme, which decomposes the likelihood over time points and observables without requiring amortization or any other adaptation of the method.
@@
Finally, applying both methods to a real dataset from a longitudinal cohort study in Ethiopia confirmed the simulation study findings and yielded broadly similar posterior approximations and epidemic trajectories. Although obtained MAP estimates differed from previously published parameter values, especially for CNFs, both methods captured the data dynamics and produced model fits with predictive energy scores close to or improving upon the deterministic analysis [Gudina et al., 2021]. By demonstrating their effectiveness in a real-world setting, our analysis underscores the potential of these methods to inform public health strategies and decision-making processes.
@@
We also compared the computational efficiency of the two methods. To ensure a meaningful comparison, both methods were calibrated to produce the same number of samples from the parameter posterior, with CNF’s neural network trained on a CPU. Despite this constraint, CNF is roughly 10 times faster than PF, due to PFs sequential nature and the higher number of simulator calls. The bulk of CNF’s computational time is spent on training, but once trained, inference can be performed in mere seconds. CNF’s amortization allows multiple inferences, providing, in this scenario, a significant advantage over PF, which requires separate computations for each run. To achieve greater speed, CNF can make use of a parallelized simulator or training the neural network on a GPU. Meanwhile, PF could benefit from the use of more advanced filtering algorithms such as guided PFs [Snyder, 2011; Chopin and Papaspiliopoulos, 2020] or using correlated particles [Wiqvist et al., 2021] to avoid degeneracy or too high particle numbers. Switching to a more advanced outer MCMC sampler such as a Gibbs sampler [Chopin and Singh, 2015] or a Metropolis-Adjusted Langevin algorithm [Lowe, 2023] and parallelization over particles could speed-up the inference procedure as well. Additionally, the lack of amortization renders assessment of calibration and hyperparameter tuning for PF too computationally expensive.
@@
The scope of this work leads to several limitations. The observation model employed additive Gaussian noise for clarity and comparability; overdispersed counts or serological adjustments may be better represented by negative‑binomial or beta‑binomial formulations in other settings. Our use of an Euler–Maruyama integration with positivity clipping introduces discretization error whose interaction with inference can lead to numerical instabilities near boundaries or in small populations. Algorithmic choices were guided by the ease of their applicability in praxis: We focused on a bootstrap PF within PMMH and a tuned CNF architecture with consistent hyperparameters and fixed simulation budget across our datasets and experimental setups. Alternative PF proposals (e.g., guided/auxiliary filters or tempering within SMC) and alternative SBI estimators (likelihood or ratio estimation) or more exhaustive hyperparameter-tuning could shift efficiency–accuracy trade‑offs. To avoid out‑of‑distribution failure modes, CNFs were trained per experiment rather than globally amortized across all data regimes, which tempers their wall‑clock advantage in this study; broader amortization remains attractive but demands careful conditioning on meta‑data such as time‑grid irregularity and noise scales [Radev et al., 2023; Wang et al., 2024].
@@
In the SBI literature, recent advances focus on workflows that narrow the gap between neural SBI and Monte Carlo by using learned proposals inside sequential Monte Carlo or MCMC [Papamakarios et al., 2019; Hermans et al., 2020]. Variational and amortized SMC approaches train proposal families to reduce weight degeneracy and improve effective sample sizes; conversely, particle methods can provide high‑quality samples to bootstrap or regularize neural estimators. Looking forward, future work in this context of hybrid approaches could focus on a flow trained in the spirit of neural posterior estimation that proposes parameter values to a PMMH kernel or guide latent‑state proposals. These methods could combine the tail coverage and speed of neural estimators with the asymptotic correctness and model flexibility of pseudo‑marginal schemes. Recent generative modeling advancements, i.e., flow matching and diffusion‑style transports, offer further avenues for robust posterior learning with stronger inductive biases and improved stability [Lipman et al., 2022; Draxler et al., 2024; Arruda et al., 2025]. Additionally, the use of adaptive prior schemes where the prior is updated in light of the available data, through approaches such as empirical Bayes, hierarchical modeling, or iterative refinement, may help to mitigate issues of non-identifiability while improving robustness. Such strategies may enhance performance in particularly challenging applications that use real world data where information is both sparse and noisy.
@@
Supporting Information
S1 Text: Supporting Information
S2 Results: Supplementary Results SIS-Model
S3 Results: Supplementary Results SIR-Model
S4 Results: Supplementary Results SEIR-Model with Dense Data
S5 Results: Supplementary Results SEIR-Model with Sparse Data
S6 Results: Supplementary Results Reparametrized SEIR-Model
S7 Results: Supplementary Results Comparison Full SEIR-Model with Reparametrized Model
S8 Results: Supplementary Results SEIR-Model with Real Data